# ———————————————————————————————————–

### v2

@@ II Quantum estimation and the displaced null measurement technique
In this section we give a brief overview of the quantum parameter estimation theory Hayashi (2005); Paris (2008); Tóth and Apellaniz (2014); Demkowicz-Dobrzański et al. (2020); Banaszek et al. (2013); Albarelli et al. (2020); Sidhu and Kok (2020) used in this paper, with an emphasis on asymptotic theory and the displaced-null measurement technique developed in our previous work Girotti et al. (2024). In particular, we explain why this method is asymptotically optimal, by employing the Gaussian approximation technique called local asymptotic normality Guţă and Kahn (2006); Guţă and Jencova (2007); Kahn and Guţă (2009); Gill and Guţă (2013); Guţă et al. (2008); Butucea et al. (2016); Yamagata et al. (2013); Fujiwara and Yamagata (2020, 2023). Later on, this picture will guide our intuition when dealing with the Markov estimation problem. For our purposes it suffices to discuss the case of one-dimensional parameters, and we refer to Girotti et al. (2024) for the multi-dimensional setting.
@@
Let $\rho_{\theta}\in M(\mathbb{C}^{d})$ be a family of quantum states depending smoothly on a one-dimensional parameter $\theta$. Consider a measurement described by a positive operator valued measure $\{M_{1},\dots,M_{k}\}$ and let $X$ be the measurement outcome with probability distribution $p_{\theta}(X=i)={\rm Tr}(\rho_{\theta}M_{i})$. The quantum Cramér-Rao bound (QCRB) Helstrom (1976); Holevo (2011); Belavkin (1976); Yuen and Lax (1973); Braunstein and Caves (1994); Braunstein et al. (1996) states that the variance of any unbiased estimator $\hat{\theta}=\hat{\theta}(X)$ is lower bounded as
@@
In general, the QCRB is not achievable when only a single copy of $\rho_{\theta}$ is available. However, the bound is attainable in the asymptotic limit of large number of samples by the following two step adaptive procedure Gill and Massar (2000). Given $n$ copies of $\rho_{\theta}$ one can use a small proportion of the samples (e.g. $\tilde{n}=n^{1-\epsilon}$ for a small $\epsilon>0$) to compute a preliminary (non-optimal) estimator $\tilde{\theta}_{n}$ of $\theta$; reasonable estimators will concentrate around $\theta$ such that $|\tilde{\theta}_{n}-\theta|=O(n^{-1/2+\epsilon})$ with high probability, which will be assumed throughout. In the second step, one measures the SLD operator $\mathcal{L}_{\tilde{\theta}_{n}}$ on each of the remaining copies. If $X_{1},\dots,X_{n^{\prime}}$ are the outcomes of these measurements (with $n^{\prime}=n-\tilde{n}$) then the estimator
@@
However, for certain models including that considered in this paper, measuring the SLD may not be feasible experimentally. Instead, we will use a different method called displaced-null measurement Girotti et al. (2024), which aims to estimate the parameter of a pure state models $\rho_{\theta}=|\psi_{\theta}\rangle\langle\psi_{\theta}|$ by measuring each copy in a basis that contains the vector $|\psi_{\tilde{\theta}}\rangle$ with $\tilde{\theta}$ close to the true parameter $\theta$. The Fisher information of such a measurement is known to converge to the QFI as $\tilde{\theta}$ approaches $\theta$ Pezze and Smerzi (2014); Pezzè et al. (2017); Liu et al. (2019). This suggests that the QCRB can be achieved asymptotically by using a two-step strategy similar to the SLD case: one first obtains a preliminary estimator $\tilde{\theta}_{n}$ and then measures each copy in a basis containing the vector $|\psi_{\tilde{\theta}_{n}}\rangle$. However, it turns out that this “null measurement” strategy fails due to the fact that for small deviations from $\tilde{\theta}_{n}$, the outcome probabilities depend on $(\theta-\tilde{\theta}_{n})^{2}$ and one cannot distinguish between left and right deviations from $\tilde{\theta}_{n}$, cf. Girotti et al. (2024) for the precise mathematical statement. This non-identifiability issue can be sidestepped by deliberately changing the reference parameter from $\tilde{\theta}_{n}$ to $\theta_{0}:=\tilde{\theta}_{n}-\delta_{n}$ where $\delta_{n}=n^{-1/2+3\epsilon}$, so that $\theta=\theta_{0}+(u+\tau_{n})/\sqrt{n}$ with $\tau_{n}=n^{3\epsilon}$. The choice of $\tau_{n}$ is not unique but we refer to Girotti et al. (2024) for the general requirements. Since $|\theta-\tilde{\theta}_{n}|=O(n^{-1/2+\epsilon})\ll\delta_{n}$, it means that $\theta$ lies on the right side of $\theta_{0}$ and can be unambiguously identified from the outcomes of a measurement in a basis $\{|e_{0}\rangle,\dots,|e_{d-1}\rangle\}$ such that $|e_{0}\rangle\equiv|\psi_{\theta_{0}}\rangle$. In addition, as $\theta_{0}$ approaches $\theta$ in the limit of large $n$, the displaced-null measurement exhibits the optimality properties of the “null measurements” without sharing their non-identifiability issues.
@@
where $f=4\|\dot{\psi}_{\tilde{\theta}_{n}}\|^{2}$ is the QFI at $\tilde{\theta}_{n}$. The estimator $\hat{\theta}_{n}$ is asymptotically optimal in the sense of equation (1) and asymptotically normal.
In appendix A we give more insight into this method by analysing its properties using the theory of local asymptotic normality Guţă and Kahn (2006); Guţă and Jencova (2007); Kahn and Guţă (2009); Gill and Guţă (2013); Guţă et al. (2008); Butucea et al. (2016). In a nutshell, for large $n$ the original model becomes equivalent to a Gaussian one consisting of a multi-mode coherent state whose amplitude depends linearly on the parameter, while the SLD and displaced-null strategies translated into measuring a quadrature and respectively the modes number operators of the shifted state. While this i.i.d. setup is different from the Markovian one studied in this paper, the overall asymptotic picture is similar and reader may find the i.i.d. case useful in guiding the intuition.
III Quantum Markov chains and post-processing using coherent absorbers
@@
A quantum Markov chain consists of a system interacting successively with a chain of independent ”noise units” (the input) modelling the environment. In this paper the system’s space is taken to be $\mathbb{C}^{d}$ while the ”noise units” are two dimensional systems prepared in the state $|0\rangle$ where $\{|0\rangle,|1\rangle\}$ is the standard basis in $\mathbb{C}^{2}$. We expect that the theory developed here works for general finite dimensional inputs, but we restrict here to this minimal setup which can be used to represent a discretised version of a continuous-time Markovian model with a single Bosonic field Attal and Pautrat (2006b).
At each time step the system interacts with the input unit via a unitary $U$ on $\mathbb{C}^{d}\otimes\mathbb{C}^{2}$. If the system is initially prepared in a state $|\varphi\rangle$, the joint state of system and noise units (output) after $n$ times steps is
@@
$=$ $U^{(n)}\cdot\dots\cdot U^{(2)}\cdot U^{(1)}|\varphi\otimes 0^{\otimes n}\rangle\in\mathbb{C}^{d}\otimes\left(\mathbb{C}^{2}\right)^{\otimes n}$
where $U^{(i)}$ is the unitary acting on the system and the $i$-th noise unit. By expanding the state (III) with respect to the standard product basis in the output we have
@@
where $K_{i}=\langle i|U|0\rangle$ are Kraus operators acting on $\mathbb{C}^{d}$.
From equation (III) it follows that the reduced system state of the system at time $n$ is given by
@@
where the partial trace is taken over the output noise units, and $\mathcal{T}_{*}:M(\mathbb{C}^{d})\to M(\mathbb{C}^{d})$ is the Markov transition operator (Schrödinger picture)
@@
whose dual (Heisenberg picture) will be denoted by $\mathcal{T}$.
@@
Throughout the paper we will assume that the dynamics is primitive in the sense that $\mathcal{T}_{*}$ has a unique stationary state $\rho_{\rm ss}>0$ so that $\mathcal{T}_{*}(\rho_{\rm ss})=\rho_{\rm ss}$ and it is aperiodic, i.e. the only eigenvalue of $\mathcal{T}_{*}$ with unit absolute value is $1$.
@@
Let $\theta\mapsto U_{\theta}$ be a smooth map describing how the dynamics depends on an unknown parameter $\theta$, which is assumed to belong to an open bounded interval $\Theta$ of $\mathbb{R}$. We use similar notations $|\Psi_{\theta,n}\rangle,K_{\theta,i},\mathcal{T}_{\theta}$ to denote the dependence on $\theta$ of the system-output state, Kraus operators, transition operator, etc. Two sequences of quantum statistical models indexed by time are of interest here: the system-output state $\mathcal{S}\mathcal{O}_{n}:=\{|\Psi_{\theta,n}\rangle:\theta\in\Theta\}$ defined in equation (III) and the output state $\mathcal{O}_{n}:=\{\rho^{\rm out}_{\theta,n}:\theta\in\Theta\}$ defined in equation (III). While the former is more informative than the latter and easier to analyse, we are particularly interested in estimation strategies which involve only measurements on the output, hence the importance of the model $\mathcal{O}_{n}$. The following Theorem Guţă and Kiukas (2015) shows that for primitive dynamics the QFI of both models scale linearly with $n$ with the same rate, so having access to the system does not change the asymptotic theory. To simplify the expression of the QFI rate (7) we assume the following ”gauge condition”
@@
The condition (6) means that $\sum_{j}\dot{K}_{\theta,j}^{*}K_{\theta,j}$ belongs to the subspace $\{X:{\rm Tr}(\rho_{\theta}^{\rm ss}X)=0\}\subseteq M(\mathbb{C}^{d})$ on which the resolvent $\mathcal{R}_{\theta}:=({\rm Id}-\mathcal{T}_{\theta})^{-1}$ is well defined as the Moore-Penrose inverse. The condition can be satisfied by choosing the complex phase of the Kraus operators appropriately, or equivalently the phase of the standard basis in the noise unit space $\mathbb{C}^{2}$.
@@
In the following, we will always assume that $f_{\theta}>0$ for every $\theta\in\Theta$. In general, the classical Fisher information associated to simple repeated measurements (measuring the same observable on each output unit) does not achieve the QFI rate $f_{\theta}$. However, the class of available measurements can be enlarged by unitarily ”post-processing” the output before performing a standard measurement, so that effectively one measures the original output in a rotated basis. While this shifts the difficulty from measurement to ”quantum computation”, it turns out that the post-processing can be implemented with minimal computational cost by employing the concept of a coherent absorber introduced in Stannigel et al. (2012). Indeed Godley and Guţă (2023) demonstrated that the QFI rate is achievable by combining post-processing by a coherent absorber with a simple adaptive sequential measurement scheme. Furthermore, Yang et al. (2023), argued that one can also achieve the QFI by performing simple counting measurements in the output, without the need for adaptive measurements. Our goal is to revisit this scheme and to provide a new, computationally and statistically effective estimation strategy.
III.1 Quantum postprocessing with a coherent absorber
The working of the coherent absorber is illustrated in Figure 1 b). Consider a QMC with a fixed and known unitary $U$. After interacting with the system, each output noise unit interacts with a separate $d$ dimensional system (the coherent absorber), via a unitary $V$. The system and absorber can now be regarded as a single doubled-up system which interacts with the input via the unitary $W:=VU$ on $\mathbb{C}^{d}\otimes\mathbb{C}^{d}\otimes\mathbb{C}^{2}$, where $U$ acts on first and third tensors and $V$ on second and third. The defining feature is that the system plus absorber have a pure stationary state. One can arrange this by requiring

### v3

@@ 2 Quantum estimation and the displaced null measurement technique
In this section we give a brief overview of the quantum parameter estimation theory [5, 6, 7, 8, 9, 10, 11] used in this paper, with an emphasis on asymptotic theory and the displaced-null measurement technique developed in our previous work [1]. In particular, we explain why this method is asymptotically optimal, by employing the Gaussian approximation technique called local asymptotic normality [117, 121, 118, 122, 119, 120, 123, 124, 125]. Later on, this picture will guide our intuition when dealing with the Markov estimation problem. For our purposes it suffices to discuss the case of one-dimensional parameters, and we refer to [1] for the multi-dimensional setting.
@@
Let $\rho_{\theta}\in M(\mathbb{C}^{d})$ be a family of quantum states depending smoothly on a one-dimensional parameter $\theta$. Consider a measurement described by a positive operator valued measure $\{M_{1},\dots,M_{k}\}$ and let $X$ be the measurement outcome with probability distribution $p_{\theta}(X=i)={\rm Tr}(\rho_{\theta}M_{i})$. The quantum Cramér-Rao bound (QCRB) [3, 4, 53, 126, 54, 127] states that the variance of any unbiased estimator $\hat{\theta}=\hat{\theta}(X)$ is lower bounded as
@@
In general, the QCRB is not achievable when only a single copy of $\rho_{\theta}$ is available. However, the bound is attainable in the asymptotic limit of large number of samples by the following two step adaptive procedure [128]. Given $n$ copies of $\rho_{\theta}$ one can use a small proportion of the samples (e.g. $\tilde{n}=n^{1-\epsilon}$ for a small $\epsilon>0$) to compute a preliminary (non-optimal) estimator $\tilde{\theta}_{n}$ of $\theta$; reasonable estimators will concentrate around $\theta$ such that $|\tilde{\theta}_{n}-\theta|=O(n^{-1/2+\epsilon})$ with high probability, which will be assumed throughout. In the second step, one measures the SLD operator $\mathcal{L}_{\tilde{\theta}_{n}}$ on each of the remaining copies. If $X_{1},\dots,X_{n^{\prime}}$ are the outcomes of these measurements (with $n^{\prime}=n-\tilde{n}$) then the estimator
@@
However, for certain models including that considered in this paper, measuring the SLD may not be feasible experimentally. Instead, we will use a different method called displaced-null measurement [1], which aims to estimate the parameter of a pure state models $\rho_{\theta}=|\psi_{\theta}\rangle\langle\psi_{\theta}|$ by measuring each copy in a basis that contains the vector $|\psi_{\tilde{\theta}}\rangle$ with $\tilde{\theta}$ close to the true parameter $\theta$. The Fisher information of such a measurement is known to converge to the QFI as $\tilde{\theta}$ approaches $\theta$ [129, 130, 131]. This suggests that the QCRB can be achieved asymptotically by using a two-step strategy similar to the SLD case: one first obtains a preliminary estimator $\tilde{\theta}_{n}$ and then measures each copy in a basis containing the vector $|\psi_{\tilde{\theta}_{n}}\rangle$. However, it turns out that this “null measurement” strategy fails due to the fact that for small deviations from $\tilde{\theta}_{n}$, the outcome probabilities depend on $(\theta-\tilde{\theta}_{n})^{2}$ and one cannot distinguish between left and right deviations from $\tilde{\theta}_{n}$, cf. [1] for the precise mathematical statement. This non-identifiability issue can be sidestepped by deliberately changing the reference parameter from $\tilde{\theta}_{n}$ to $\theta_{0}:=\tilde{\theta}_{n}-\delta_{n}$ where $\delta_{n}=n^{-1/2+3\epsilon}$, so that $\theta=\theta_{0}+(u+\tau_{n})/\sqrt{n}$ with $\tau_{n}=n^{3\epsilon}$. The choice of $\tau_{n}$ is not unique but we refer to [1] for the general requirements. Since $|\theta-\tilde{\theta}_{n}|=O(n^{-1/2+\epsilon})\ll\delta_{n}$, it means that $\theta$ lies on the right side of $\theta_{0}$ and can be unambiguously identified from the outcomes of a measurement in a basis $\{|e_{0}\rangle,\dots,|e_{d-1}\rangle\}$ such that $|e_{0}\rangle\equiv|\psi_{\theta_{0}}\rangle$. In addition, as $\theta_{0}$ approaches $\theta$ in the limit of large $n$, the displaced-null measurement exhibits the optimality properties of the “null measurements” without sharing their non-identifiability issues.
@@
where $f=4\|\dot{\psi}_{\tilde{\theta}_{n}}\|^{2}$ is the QFI at $\tilde{\theta}_{n}$. The expression of $\hat{u}_{n}$ is derived in section 7 by taking into account that the pattern counts are asymptotically normal (conditional to the preliminary estimator $\tilde{\theta}_{n}$) and finding the unbiased linear combination with the smallest variance, which turns out to be related to the total number of patterns as above. The estimator $\hat{\theta}_{n}$ is asymptotically optimal in the sense of equation (1) and asymptotically normal.
In appendix A we give more insight into this method by analysing its properties using the theory of local asymptotic normality [117, 121, 118, 122, 119, 120]. In a nutshell, for large $n$ the original model becomes equivalent to a Gaussian one consisting of a multi-mode coherent state whose amplitude depends linearly on the parameter, while the SLD and displaced-null strategies translated into measuring a quadrature and respectively the modes number operators of the shifted state. While this i.i.d. setup is different from the Markovian one studied in this paper, the overall asymptotic picture is similar and reader may find the i.i.d. case useful in guiding the intuition.
3 Quantum Markov chains and post-processing using coherent absorbers
@@
A quantum Markov chain consists of a system interacting successively with a chain of independent ”noise units” (the input) modelling the environment. In this paper the system’s space is taken to be $\mathcal{H}_{s}\cong\mathbb{C}^{d}$ while the ”noise units” are two dimensional systems prepared in the state $|0\rangle$ where $\{|0\rangle,|1\rangle\}$ is the standard basis in $\mathcal{H}_{n}\cong\mathbb{C}^{2}$. We expect that the theory developed here works for general finite dimensional inputs, but we restrict here to this minimal setup which can be used to represent a discretised version of a continuous-time Markovian model with a single Bosonic field [132].
At each time step the system interacts with the input unit via a unitary $U$ on $\mathcal{H}_{s}\otimes\mathcal{H}_{n}$. If the system is initially prepared in a state $|\varphi\rangle$, the joint state of system and noise units (output) after $n$ times steps is
@@
$=$ $U^{(n)}\cdot\dots\cdot U^{(2)}\cdot U^{(1)}|\varphi\otimes 0^{\otimes n}\rangle\in\mathcal{H}_{s}\otimes\left(\mathcal{H}_{n}\right)^{\otimes n}$
where $U^{(i)}$ is the unitary acting on the system and the $i$-th noise unit. By expanding the state (3) with respect to the standard product basis in the output we have
@@
where $K_{i}=\langle i|U|0\rangle$ are Kraus operators acting on $\mathcal{H}_{s}$.
From equation (3) it follows that the reduced system state of the system at time $n$ is given by
@@
where the partial trace is taken over the output noise units, and $\mathcal{T}_{*}:L_{1}(\mathcal{H}_{s})\to L_{1}(\mathcal{H}_{s})$ is the Markov transition operator (Schrödinger picture)
@@
whose dual (Heisenberg picture) will be denoted by $\mathcal{T}$. Here $L_{1}(\mathcal{H}_{s})$ denotes the trace-class operators on $\mathcal{H}_{s}$, which in this case is isomorphic to $M(\mathbb{C}^{d})$.
@@
Throughout the paper we will assume that the dynamics is primitive in the sense that $\mathcal{T}_{*}$ has a unique stationary state $\rho^{\rm ss}>0$ so that $\mathcal{T}_{*}(\rho^{\rm ss})=\rho^{\rm ss}$ and it is aperiodic, i.e. the only eigenvalue of $\mathcal{T}_{*}$ with unit absolute value is $1$.
In order to make the presentation of our results more accessible, we will show how the theory we develop applies to a simple example; we will consider a two dimensional system whose reduced dynamics (precession around $z$ axis with bit-flip noise) is given by the following Lindblad evolution:
$\frac{d\rho_{t}}{dt}=-i\omega[\sigma_{z},\rho_{t}]+\gamma(\sigma_{x}\rho_{t}\sigma_{x}-\rho_{t}),$ (6)
where $\omega,\gamma>0$ and $\sigma_{x}$, $\sigma_{z}$ is the usual notation for $z$ and $x$ Pauli matrices. Both the frequency $\omega$ or the coupling strength $\gamma$ are natural parameters to estimate. In the input-output formalism, the joint system and output state is given by a continuous matrix product state; considering a small enough time discretization $dt$, one can approximate it by a state of the form as in equation (4) where the Kraus operators are given by
$K_{0}=\sqrt{1-\gamma_{t}}e^{-i\omega_{t}\sigma_{z}},\quad K_{1}=e^{-i\omega_{t}\sigma_{z}}\sqrt{\gamma_{t}}\sigma_{x},$
$\gamma_{t}=\gamma dt$, $\omega_{t}=\omega dt$ Both the continuous and the discrete time evolutions have the maximally mixed state $\mathbf{1}/2$ as unique stationary state.
@@
Let $\theta\mapsto U_{\theta}$ be a smooth map describing how the dynamics depends on an unknown parameter $\theta$, which is assumed to belong to an open bounded interval $\Theta$ of $\mathbb{R}$. We use similar notations $|\Psi_{\theta,n}\rangle,K_{\theta,i},\mathcal{T}_{\theta}$ to denote the dependence on $\theta$ of the system-output state, Kraus operators, transition operator, etc. Two sequences of quantum statistical models indexed by time are of interest here: the system-output state $\mathcal{S}\mathcal{O}_{n}:=\{|\Psi_{\theta,n}\rangle:\theta\in\Theta\}$ defined in equation (3) and the output state $\mathcal{O}_{n}:=\{\rho^{\rm out}_{\theta,n}:\theta\in\Theta\}$ defined in equation (3). While the former is more informative than the latter and easier to analyse, we are particularly interested in estimation strategies which involve only measurements on the output, hence the importance of the model $\mathcal{O}_{n}$. The following Theorem [113] shows that for primitive dynamics the QFI of both models scale linearly with $n$ with the same rate, so having access to the system does not change the asymptotic theory. To simplify the expression of the QFI rate (8) we assume the following ”gauge condition”
@@
The condition (7) means that $\sum_{j}\dot{K}_{\theta,j}^{*}K_{\theta,j}$ belongs to the subspace $\{X:{\rm Tr}(\rho_{\theta}^{\rm ss}X)=0\}\subseteq B(\mathcal{H}_{s})$ of bounded operators, on which the resolvent $\mathcal{R}_{\theta}:=({\rm Id}-\mathcal{T}_{\theta})^{-1}$ is well defined as the Moore-Penrose inverse. The condition can be satisfied by choosing the complex phase of the Kraus operators appropriately, or equivalently the phase of the standard basis in the noise unit space $\mathcal{H}_{n}$.
@@
In the following, we will always assume that $f_{\theta}>0$ for every $\theta\in\Theta$. In general, the classical Fisher information associated to simple repeated measurements (measuring the same observable on each output unit) does not achieve the QFI rate $f_{\theta}$. However, the class of available measurements can be enlarged by unitarily ”post-processing” the output before performing a standard measurement, so that effectively one measures the original output in a rotated basis. While this shifts the difficulty from measurement to ”quantum computation”, it turns out that the post-processing can be implemented with minimal computational cost by employing the concept of a coherent absorber introduced in [116]. Indeed [115] demonstrated that the QFI rate is achievable by combining post-processing by a coherent absorber with a simple adaptive sequential measurement scheme. Furthermore, [2], argued that one can also achieve the QFI by performing simple counting measurements in the output, without the need for adaptive measurements. Our goal is to revisit this scheme and to provide a new, computationally and statistically effective estimation strategy.
3.1 Quantum postprocessing with a coherent absorber
The working of the coherent absorber is illustrated in Figure 1 b). Consider a QMC with a fixed and known unitary $U$. After interacting with the system, each output noise unit interacts with a separate $d$ dimensional system $\mathcal{H}_{a}\cong\mathbb{C}^{d}$ (the coherent absorber), via a unitary $V$. The system and absorber can now be regarded as a single doubled-up system which interacts with the input via the unitary $W:=VU$ on $\mathcal{H}_{s}\otimes\mathcal{H}_{a}\otimes\mathcal{H}_{n}$, where $U$ acts on first and third tensors and $V$ on second and third. The defining feature is that the system plus absorber have a pure stationary state. One can arrange this by requiring